OpenStudy (thomas5267):

Is it possible to prove that \(a_1+b_1i=a_2+b_2i\) if and only if \(a_1=a_2\) and \(b_1=b_2\) or is it a definition?

OpenStudy (ikram002p):

i ws sure that its a property , but wanted tomake sure from my txt book this is exactly what it said :- http://prntscr.com/4551o2

OpenStudy (thomas5267):

So it is a definition and it is impossible to prove that it is true?

OpenStudy (ikram002p):

wait i missed something hehe since its a properties , that means its axiom sys assumtion :) and its not def , its prop and this is what i ment to post ealiear http://prntscr.com/4553c8

OpenStudy (ikram002p):

there is a deffrence btw defenition and a proparty

OpenStudy (thomas5267):

How do you define equality? I looked up the axioms of field on Wikipedia but the axioms assume a equality relation.

OpenStudy (ikram002p):

equality is definition , also imaginary part and real part but saying iff reals part are equal , and imaginary part are equal is a property :D

OpenStudy (ikram002p):

got it nw ? or still confused ? i can explain axiom system to you if you want :)

terenzreignz (terenzreignz):

Maybe it is? haha Suppose \[\Large a_1 + \color{blue}ib_1 = a_2 +\color{blue} ib_2\]

terenzreignz (terenzreignz):

and I don't know what I'm doing lol

OpenStudy (ikram002p):

lolz tere , well what happen is this 1_ we assumed that this statment is true in complex plane system " that a1+b1i=a2+b2i if and only if a1=a2 and b1=b2 " ( property) 2_and we can use this in other proofs in the same system , how ever this property it self is stick xD no way to prove or disprove it

terenzreignz (terenzreignz):

Okay... idea :D The above assumption implies that \[\Large (a_1 - a_2) + \color{blue}i(b_1 - b_2) = 0\]

OpenStudy (ikram002p):

terez lol ur going on circle proof :P

terenzreignz (terenzreignz):

Hear me out first...lol

terenzreignz (terenzreignz):

We can then can take the modulus of both sides... and they should still be equal, obviously. \[\Large (a_1- a_2)^2 + (b_1+b_2)^2 = 0\]

OpenStudy (ikram002p):

ok lol im all ears :)

terenzreignz (terenzreignz):

And squares of real numbers are nonnegative.

terenzreignz (terenzreignz):

And yet we have \[\Large (a_1 - a_2)^2 = -(b_1 - b_2)^2\]

terenzreignz (terenzreignz):

We can say with confidence that the left-side is nonnegative. We can say with confidence that the right-side is nonpositive.

terenzreignz (terenzreignz):

That can only be true if \[\Large a_1 - a_2 = 0 = b_1 - b_2\]

terenzreignz (terenzreignz):

Sound good? ^_^

OpenStudy (ikram002p):

haha and ?

terenzreignz (terenzreignz):

Well, \[\Large a_1 - a_2 =0 \qquad \rightarrow \qquad a_1 = a_2\] And same goes for the b's XD

OpenStudy (ikram002p):

mmm cool xD

ganeshie8 (ganeshie8):

next, thomas is going to ask us to prove : \(\large z_1=z_2 \implies |z_1| = |z_2|\)

terenzreignz (terenzreignz):

Heaven help us D:

OpenStudy (thomas5267):

Can we define equality in the first place? If not, this is a wild goose chase.

terenzreignz (terenzreignz):

Oh noes... philosophy... Nothing to do here haha

OpenStudy (ikram002p):

wow , ur going into equality definition which is undefined term , right @ganeshie8

ganeshie8 (ganeshie8):

i never took analysis before, so im bit clueless on where exactly im going lol

OpenStudy (thomas5267):

"In modern first-order logic equality is primitive. We don't define it, we just know that two things are equal if and only if they are the same thing. So two numbers are equal if and only if they are the same number." http://math.stackexchange.com/questions/855903/how-do-we-define-equality-in-real-numbers

OpenStudy (ikram002p):

so undefined term

OpenStudy (thomas5267):

Okay. Assuming that equality on real numbers is defined, can you derive the equality in complex number?

OpenStudy (ikram002p):

me ? , no :P i stick with axioms and stay in the safe side

OpenStudy (thomas5267):

It seems like we need Zermelo–Fraenkel set theory to solve this simple question... ._.

OpenStudy (ikram002p):

yep , so you want it ? i think i have it in my note

OpenStudy (ikram002p):

i guess all complex reasearch / books mintion that operation are the same in real or exaclty they say The comple number system is , a natural extention of real number system

OpenStudy (thomas5267):

This question looks ridiculously complex now. First we have to use ZFC. Then we have to define real numbers. Thirdly, we can now use something called Cayley–wingspanson construction to construct the complex. Finally, we can (hopefully) derive this definition of equality. Too much work... I give up...

OpenStudy (thomas5267):

OpenStudy is dumb. D!ckson becomes wingspanson...

OpenStudy (ikram002p):

haha lolz u dnt have to do this work , its already exist only small lovely paper of axioms on complex u need :)

OpenStudy (thomas5267):

Where is that small and lovely piece of paper?